OpenStudy (snopet):

I just took a geometry tests on proofs and got this question wrong: Adam's proof: angle 1 + angle 2 + angle 3 + angle 4 = 360° Therefore, angle 2 + angle 3 = 180° (t is a straight line) Hence, angle 1 = angle 3 (Transitive Property of Equality) Darius' proof: angle 1 + angle 4 = 180° (t is a straight line) angle 1 + angle 2 = 180° (PQ is a straight line) Therefore, angle 1 + angle 2 = angle 1 + angle 4 (Transitive Property of Equality) Hence, angle 2 = angle 4 (Subtraction Property of Equality)

OpenStudy (snopet):

These were the answer choices: Only Adam's proof is correct. Both Adam's and Darius' proofs are correct. Both Adam's and Darius' proofs are incorrect. Only Darius' proof is correct. I chose that both were incorrect even though Darius' broth was correct because he said that "T was is a straight line." Shouldn't it be because they are supplementary angles?

OpenStudy (snopet):

(Please ignore my typos.)

likeabossssssss (likeabossssssss):

i srry i dont rember but may be @Awolflover1 could help chu <3

likeabossssssss (likeabossssssss):

@AloneS

OpenStudy (mhchen):

I'm guessing that they were perpendicular bisectors. Because angle 1 and angle 4 are supplementary. angle 1 and angle 2 are supplementary  If 3 angles are next to each other, and each is supplmentary of each, then they are all 90 degrees.

iYuko (iyuko):

that is accurate

OpenStudy (mhchen):

That's graphically. Now based on the proof: angle 1 and angle 4 are supplementary angle 1 and angle 2 are supplementary And hence 1 + 2 = 1 + 4. That way, 2 = 4

OpenStudy (snopet):

So I was correct in stating that neither proof was right?

OpenStudy (mhchen):

When it says 't' is a straight line, did it have a graph?

OpenStudy (snopet):

I FORGOT TO POST THE PHOTO I'M SORRY!

OpenStudy (mhchen):

Depending on if there was a diagram provided showing that angle 1 and angle 4 are on a straight line.

OpenStudy (mhchen):

Okay.....well you're right that it should be supplementary angles. I guess your teacher made the assumption that straight lines always give supplementary angles. :\ I would argue that to your teacher though to see if you can get a point back.

OpenStudy (mhchen):

Like seriously, that's what I would do. (Unless you guys have it in notes that straight line gives supplementary angles, then you're kinda screwed lel)

OpenStudy (snopet):

I did she told me that "sometimes you have to learn what's important and what's not" Not sure what that means... Thank you anyways!

OpenStudy (snopet):

she's gotten answers wrong before and wouldn't give me the points... Oh well :/

OpenStudy (mhchen):

wow......it's those kind of teachers xD

OpenStudy (snopet):

Yup! Fun!